### v1

# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

Abstract.
In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk with the user during a collaborative assembly task, and explore how the presence of negative team feedback may affect team dynamics and blame attribution. Through an exploratory study with 20 participants, we found that 18 individuals interacted socially with the robot, nine of which initiated questions back to the robot. We report the frequency and length of users’ responses in task-oriented and non-task-oriented dialogue, and further elaborate on people’s reactions to the negative system feedback and robot-initiated small talk. We discuss the potential for integrating small talk in non-social robots, and propose three design guidelines to enhance human-robot small talk interactions.
Human-Robot Collaboration, Social Robots, Small Talk, HRI
ccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
This figure includes 6 photos, alternating between the participant themselves and the participant and the robot interacting together. Below is text for each image that depicts the conversation that is exchanged between the human and the robot. The conversation is as follows: robot: on another note, do you like to play any games?; participant: um, I like Tetris if you if you’ve ever played that?; participant: Can I get another yellow pipe?; robot: Getting it now.; robot: No, I have not; participant: you might like it; robot: Can you tell me more about it?; participant: [chuckle] I think it’s a pretty classic game you know you just like the pieces different shape pieces are dropping down and you’re trying to get them to fill up a line across and then it’ll like clear it. they even made a movie about it.
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
1. Introduction
Human-robot collaboration research envisions robots as human collaborators, or co-workers, to enhance human work in diverse domains. Take manufacturing as an example: the manufacturing industry is currently undergoing a significant transformation to become more flexible and agile, allowing it to accommodate mass product customization even while product life cycles are continuously decreasing. Moreover, small and medium-sized manufacturing companies face a shortage of skilled production workers. Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable flexible manufacturing, and increase productivity.
Though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks. Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (kopp2021success, 21). While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé_Mutlu_2015, 41), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor. It is unclear if people would engage in meaningful social interactions with non-anthropomorphic robot manipulators.
In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver1981LinguisticRA, 24)—as human co-workers would. Specifically, we sought to study whether and how people respond to small talk from a robot manipulator during a collaborative task. Moreover, we were interested in understanding how negative performance feedback to the human-robot team might shape the team dynamics and blame attribution; this inquiry is relevant and of interest because of the rising adoption of performance monitoring systems in industrial settings (e.g., warehouse). Through an exploratory study with 20 participants, we found that most participants (18) interacted socially with the robot (Fig. 1), nine of which even initiated questions back to the robot. Overall, the robot’s small talk led to pleasant experiences and did not negatively impact task efficiency. We discuss the potential for integrating small talk in collaborative robots, and present design guidelines to enhance human-robot small talk interactions.
2. Background and Related Work
2.1. Social Human-Robot Interactions
There have been efforts to study how robots can interact socially with humans in various domains such as for tutoring and educational contexts (belpaeme2018education, 5) (salomons2022we, 40), encouraging exercise (antony2023co, 2), providing companionship (kidd2006sociable, 18) (tsoi2021challenges, 50) and emotional support (sabelli2011conversational, 39), and other therapeutic interventions (feil2009toward, 13)(scassellati2018improving, 42). Many of these engagements explore the integration of various multi-modal capabilities in robots to promote their engagements with humans (youssef2022survey, 53). These can consist of robot eye gaze behaviors (mutlu2009footing, 32) (admoni2017social, 1), head movements (sidner2006effect, 45) (kkedzierski2013emys, 17) and gestures (meena2012integration, 30) alongside robot speech abilities (striepe2021modeling, 48) (sinnema2019attitude, 46). While research has been conducted to analyze the characteristics of conversational interactions with robots (seok2022cultural, 44) (laban2020tell, 22), the majority of these works utilize social robots as their robotic entity. However, the prevalence of non-social robots such as cobots or other manipulator robots is much greater than social robots themselves, and few efforts have been made to further explore what will happen if we have conversations between humans and non-social robots (chowdhury2021you, 8) (li2022tod, 26).
2.2. People Treat Machines Like People
The inclination of humans to treat machines socially can be contextualized within broader theoretical frameworks. The Computers Are Social Actors (CASA) paradigm (nass1994computers, 33) and the Media Equation Theory (reeves1996media, 36) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners. Such anthropomorphic tendencies highlight the intersection of technology and human social instincts. Even in industrial manufacturing environments, there remains an intrinsic human desire for social interaction. As operators often partake in casual conversations or “small talk” during shifts (cheon2022robots, 7), they have also found themselves speaking to the robot they work with (Sauppé_Mutlu_2015, 41). This suggests that individuals not only expect but actively desire social features in manufacturing robots.
Some studies explored the potential of enhancing cobots primarily designed for physical manipulation tasks with more human-like movements (hostettler2023human, 16). One study specifically integrated principles from character animations (terziouglu2020designing, 49) to enhance the perception of cobots as social entities. Such amalgam of technology and socialization emphasizes the importance of recognizing and harnessing the persistent human drive for social connection, even with machines. Given the limited research on designing social behaviors for industrial robots and manipulators with a non-human form, despite its potential, a pressing question arises: How can we further design human-like attributes for robots that are inherently devoid of such characteristics? Building on this notion, our exploratory study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
2.3. Small Talk
Social dialogue that consists of light conversation between new or unfamiliar parties is often referred to as “small talk” (Laver1981LinguisticRA, 24). Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (coupland2003small, 10, 35). It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (holmes2005small, 15). While prevalent in American culture, attitudes towards the value of small talk differ across cultures (bickmore2005social, 6, 38, 52).
The integration of small talk to non-human agents engaging in conversations with people has been explored to help promote trust and rapport between the human and the agent. Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (clark2019good, 9, 14, 20), voice assistants (mahmood2023llm, 28), and virtual agents (zhou2019trusting, 54). Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (babel2021small, 4), facial expressions (paradeda2016facial, 34), or emotional displays (aroyo2018trust, 3), has been shown to elevate user trust levels. Furthermore, researchers have investigated which behaviors in particular can build rapport between humans and a collaborative robot (seo2018investigating, 43). However, less explored is the domain of applying small talk to foster rapport between humans and other agents with physical capabilities that are not typically poised to engage in conversations with people, such as autonomous self driving vehicles (large2019s, 23), or industrial robots (li2023irwoz, 25). Preliminary studies hint at the benefits of integrating small talk with task-oriented dialogue in industrial robot settings (li2022tod, 26). Yet, understanding users’ interactions and experiences with such robots remains unexplored. Investigating individuals’ responses and reactions to small talk initiated by non-humanoid robots is essential for comprehending its influence on human-robot collaborations and gathering insights on how to foster better human-robot partnership.
3. METHODS
We conducted an exploratory study to analyze people’s behaviors and attitudes towards an industrial robot that exhibited social behavior (via initiated small talk) while examining the change in human-robot team dynamics after the team receives team feedback.
3.1. Collaborative Assembly Task
Participants were informed that the assembly task was intended to mimic a real-world collaboration setup between a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace. They were given a fictional team assignment (Team 3) and were told the collaborative task scenario was a prototype for a larger-scale warehouse interaction. It was intended to simulate a modern warehouse that may contain a computerized efficiency monitoring system to oversee the team collaboration. The monitoring system would communicate with them via speech but would be distinguishable from any speech communication coming from the robot.
Each participant engaged in a collaborative PVC pipe structure building task with a robot manipulator mounted on a table. They were given diagrams corresponding to the structures they were required to assemble. The participants read instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe. The robot would provide a verbal confirmation message if a request was received correctly and an additional instruction reminder message once per structure when the participant requested their second to last pipe for that structure. The pipe would be placed on a ramp that connected the robot’s workspace to the participant’s workspace so that the pipe could be delivered to the participant without the robot entering their workspace. Participants were provided with joints in containers that corresponded to each structure to be built. To simulate a realistic industrial collaborative environment that might involve the use of defective materials, each participant was given one faulty joint for every pipe structure. They were instructed to inspect each joint before using it in the structure. If participants identified a faulty joint (marked with red sticker), they would place the joint on a ramp that would deliver it to a container inside the robot’s workspace. The robot would retrieve the joint when given a verbal command such as ”Report a faulty joint”. The experimental setup is shown in Figure 2.
Figure 2. The experimental setup. The robot can grab a pipe from the pipe dispenser and place it on a pipe ramp, which delivers it to the participant. The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
The experimental setup. The top image shows two tables next to one another. The table on the left is the participant workspace and contains four boxes with pipe connectors that the participant can grab joints from. The table on the right is where the robot is mounted and serves as the robot’s workspace. A ramp that the participant can use for the delivery of faulty pipes to the robot and a ramp that the robot can use to deliver requested pipes to the participant connect the two tables. The bottom image shows another view of the two tables, with a closer view of the robot workspace that highlights a pipe dispenser to the right of the robot from which the robot collects the pipe the participant requested.
Figure 2. The experimental setup. The robot can grab a pipe from the pipe dispenser and place it on a pipe ramp, which delivers it to the participant. The participant can grab a joint from the joint containers on the left and place a faulty joint on the joint ramp for it to be retrievable by the robot.
3.2. System Manipulation and Robot Behavior
We present the system and robot behavior manipulations implemented in our exploratory study to understand how the social facets of a robot manipulator influence users’ perceptions of the robot.
3.2.1. Negative Team Feedback
Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system. Each message was prefaced by an audible beep:
  1. (1)
    Feedback 1: “System announcement. Team 3, please work on improving your performance.”
  2. (2)
    Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
3.2.2. Social Behavior Robot Small-Talk
The robot exhibited four types of social and task-oriented speech behaviors: questions, statements, reminders and confirmations. During the main assembly task, the robot would engage in small talk by initiating questions and/or statements after it delivers a request confirmation message or an instruction reminder message. The robot initiated a minimum of one question or statement per request, regardless of the response rate of the participants. After a break in conversation via the negative feedback announcement, the robot initiated questions and statements on a different topic. The conversation touched on a range of questions spanning three primary categories to accommodate the diverse interests of individuals: travel; books, TV shows, and movies; and sports and games. Some examples of robot small-talk phrases are shown in Table 1.
Table 1. Examples of robot questions and statements
Topic Phrase
Travel By the way, do you like to travel?
I arrived to the U.S. from Germany in a large box.
Books, TV, Movies What do you like about those books?
That sounds like a great way to relax.
Sports, Games Do you have a favorite sport to watch?
It is good to spend time outdoors exercising.
3.3. System Implementation
3.3.1. Robot System
We utilized a Panda robot from Franka Emika. It is an industrial robotic arm with 7 degrees of freedom (7-DoF). The robot was restricted to its designated workspace and never intruded into the participant’s allocated area. Its movements were preprogrammed to execute specific sequences of actions. Among these, two involved pick-and-place motions: one to retrieve and deliver either a green or yellow PVC pipe to the participant, and the other to pick up a container holding a PVC pipe joint and subsequently move the container over a bin to discard the joint. The robot interacted with participants using text-to-speech communication.
3.3.2. Wizard-of-Oz Control
The experimenter controlled the robot using a Wizard-of-Oz (WoZ) approach (woz2012riek, 37). Through a remote graphical user interface (GUI), the experimenter executed robot actions and selected from preprogrammed conversational speech phrases. Robot actions and initial task messages were triggered using one of five GUI buttons. For conversational speech, the experimenter inputted specific alphanumeric codes into a textbox. Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via the text-to-speech feature using Google Text-to-Speech (TTS) online.
3.3.3. Preprogrammed Speech Behavior
When a robot movement was triggered, a confirmation message was automatically activated to notify the participant that the action would occur. Robot confirmation messages were preprogrammed and selected randomly during each execution. Task reminder messages were similarly preprogrammed and triggered automatically when the participant had requested enough pipes to be one away from completing a structure. Negative feedback directed towards the human-robot team was triggered automatically when participants had requested half the pipes needed for the first and second structure.
3.4. Study Procedure
Participants first completed the consent form and a demographics questionnaire. Afterward, the experimenter explained the task scenario and gave the participant instructions about the task and how to communicate with the robot. After the experimenter left the room, the participant received an introductory message from the monitoring system. The participant then engaged in a practice session to construct one pipe structure, allowing them to become acquainted with the robot’s task-oriented behaviors and the use of verbal commands. The robot only exhibited task-oriented dialogue during the practice task, such as a reminder message to move their pipe structure over to another table when finished and a confirmation message signifying it comprehended the request. When finished, the participant informed the experimenter that they were done. Following this, they took part in a main session where they built three additional pipe structures with the robot. After completing the structures, the participant informed the experimenter that they were done. At the end of the study, the experimenter conducted a semi-structured interview with each participant to understand their experience collaborating with the robot.
3.5. Participants
24 participants were recruited through convenience sampling, out of which 4 were excluded due to technical issues. The remaining 20 participants (11 female, 9 male) ranged in age from 18 to 38 ($M=24.35,SD=5.44$). Participants had a variety of educational backgrounds, including computer science, engineering, natural sciences, education, biotechnology, administration and international studies. Twelve participants listed Asian as their ethnicity; five as Caucasian; one as Hispanic, Latino, or Spanish origin of any race; one as Asian and Caucasian; and one as Caucasian and Hispanic, Latino, or Spanish origin of any race. Participants were familiar with technology ($M=4.15,SD=0.75$, 5-point scale with 5 being Very Familiar and 1 being Not At All), had moderate experience with robots ($M=2.5,SD=1$), and had little experience teaming with robots ($M=1.65,SD=0.93$). The study took about 45 minutes and was approved by the university’s institutional review board. Participants were compensated with \$15 US dollars.
3.6. Data Analysis
We collected audio-visual data of the participants’ collaboration with the robot during the pipe assembly tasks, and of the semi-structured interview. Each participant had an average of 12 minutes of interaction with the robot over the main tasks. The experimenter transcribed the audio data and conducted thematic analysis on the transcripts to identify central themes and attitudes surrounding the participants’ interactions with the robot. The participants’ verbal phrases during the main session were coded by the context in which it was made (i.e., request, response to robot task-oriented message, response to robot small-talk, question initiated to the robot). The participant small talk responses were further coded to identify the general length of the response: vocable, short sentence (¡= 4 words), medium-length sentence (¿ 4 words), and multiple sentences. We deduced response rates based on code frequencies from the interaction data. Response rates for the robot’s questions, statements, reminders, and confirmations were calculated as the percentage of those prompts that received a user response.
Given the robot’s non-human-like appearance, people might find its small talk unexpected. Thus, to evaluate the impact of small talk in robots primarily designed for physical collaboration, we analyzed the participants’ reactions to the robot’s initial small talk conversational attempt via three different aspects:
1. Conversation reciprocation: We coded a person’s response to the robot’s first conversational attempt as no response or response, and also looked at if the response was an answer to a robot question.
2. Impact on task: To observe whether the robot’s question disrupted the participant’s ongoing task, we code people’s behavior in response to the robot’s first small talk question into three actions: minimal (quick glance at the robot but no task disruption), freeze (stops current action and looks at the robot before continuing the task), and none (did not stop or look at the robot).
3. Affective reactions: To gather insights into the participants’ emotional or mental state, we grouped people’s non-verbal reactions during the robot’s first question into three categories: pleasant (smile), surprise (chin raise, lower lip raise, brow lowering etc.), and indifference (no reaction).
4. Results
In this exploratory study, first, we aim to understand whether, how, and to what extent participants engage with robot-initiated small talk. We examine the response rate, length of responses, and user-initiated questions. Second, we assess how participants perceive the robot’s initial small talk question, specifically its impact on their ongoing action and emotional state. Third, we investigate whether a negative evaluation of the human-robot team influences participants’ perceptions of the robot and their teamwork; this involves analyzing their verbal and non-verbal reactions to negative feedback and determining whom they blame for such feedback.
4.1. Participant Response Rates
Of the 20 participants, two participants did not engage in additional conversational interaction with the robot apart from the verbal commands needed to complete the collaborative task. We determined the question response rate for all participants to be 86.67% (325 out of 375), the statement response rate as $36.27\%$ (70 out of 193), and the task-oriented response rate for confirmation and reminder messages as 1.68% (8 out of 475) and 16.67% (10 out of 60), respectively. For the 18 participants who did engage in small talk, we calculated the question response rate to be 91.55% (325 out of 355), the statement response rate as $39.11\%$ (70 out of 179), and the task-oriented response rate for confirmation and reminder messages as 1.87% (8 out of 427) and 18.52% (10 out of 54), respectively.
4.2. How Participants Respond to Small Talk
4.2.1. Length of participant responses
There were a total of 395 verbal responses made to the robot for non-task-oriented dialogue; examples are shown in Table 2. Participants’ responses were mostly one sentence ($N=345$, $87.34\%$). The one-sentence responses were further broken down as follows: vocable ($N=11$, $2.78\%$), short sentence (¡= 4 words; $N=168$, $42.53\%$), and medium-length sentence (¿ 4 words; $N=166$, $42.03\%$). The remaining participant responses were comprised of multiple sentences ($N=50$, $12.66\%$), reflecting higher and possibly meaningful engagement.
Table 2. Examples of People’s Responses That Vary in Length
Category Conversation
R: In the US, I have traveled from the third floor
of this building to the second floor.
Vocable P: Whaaat? (P20)
R: Do you like to play sports?
Short Sentence P: No, not really. (P6)
R: How often do you watch movies?
Medium Sentence
P: Not that often, like twice three times
a month. (P20)
R: Where have you been in the United States?
Multi-Sentence
P: Most of the East Coast, a lot of the West
Coast, the Midwest. I was born in Utah. Uh, I’ve
been to a lot of places in the United States. (P6)
4.2.2. Participant-initiated questions
Apart from answering the robot-initiated small talk questions, we observed that participants also asked the robot questions. Nine participants ($45\%$) asked a total of 46 questions to the robot. While some asked questions back as a matter of courtesy, others asked out of genuine curiosity. P1 explained, ”I wanted to know more about the robot, like once it started talking to me”.
Furthermore, the questions asked to the robot were often personal, asking it about its preferences or capabilities. P7 remarked, ”Because he’s a robot and I’m not familiar with a robot and …I …truly want to know …how long does it take to get assemble …he said, …he’s from Germany by a group, but I’m more …interested …I tried to ask …do you like your owner?” For the eleven participants that did not initiate a question to the robot, many described their perception of robot as non-human-like as a contributing factor. One participant commented how the conversation, ”just doesn’t have as much value as …if I was talking to an actual person and hearing where they want to travel”(P13). For the two participants that did not engage with the robot at all, one stated ”It’s not a human, you don’t have a conversation” (P17). The other explained how ”from beginning to end, I always felt like it’s a robot” (P3).
4.3. Participant Response to the First Robot Small Talk Question
In addition to understanding participants’ respond to robot small talk, we are interested in their first reaction to the robot’s small talk question. We present our findings on three different aspects of the their responses, see Figure 3.
4.3.1. Conversation reciprocation
18 participants responded to the robot’s first question: “By the way, do you like to travel?”. 17 participants responded with an answer similar to e.g., “Yeah, I like traveling”. However, one participant replied with a rhetorical question: “Why are you talking to me robot? This has nothing to do with the task.” (P12). Only two participants did not respond to the robot’s question; the same two participants who did not engage in any non-task conversation with the robot.
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
Figure 3: A figure graphically displaying the participant’s response to the first robot small task question (e.g., ”By the way, do you like to travel?”) in terms of conversation reciprocation, impact on task, and affective reaction, with numbers listed for each type of participant response (conversation reciprocation: 17 answer, 2 no engagement, 1 refuse; impact on task: 13 minimal [quick glance], 4 freeze, 3 none [no behavioral acknowledgment]; affective reaction: 6 indifference, 6 surprise, 6 smile)
Figure 3. Different aspects of participants’ response to the first robot small talk initiation: (1) conversational reciprocation, (2) impact on task, and (3) affective reaction. For each aspect, the states are mutually exclusive. The numbers denote participants who exhibited the corresponding responses.
4.3.2. Impact on task
The majority of participants ($N=13$) merely exchanged a brief glance with the robot when it posed the first question; they looked at it momentarily, without letting it interrupt their ongoing task, suggesting that their attention was momentarily captured by the robot’s ”unusual” behavior (initiating small talk), but not to a degree that proved disruptive. In contrast, three participants did not look at the robot at all. Interestingly, four participants momentarily halted their task for at least two seconds, looked intently at the robot (indicating a ”freeze” reaction), and then resumed their activity. These individuals also engaged in a verbal response to the robot.
4.3.3. Affective reaction
Eight out of the 20 participants smiled in response to the robot’s first question, indicating a pleasant emotional state. Six participants displayed signs of surprise through various non-verbal cues such as chin raiser, lower lip raiser, and eye-brow lowering. The remaining six participants exhibited indifference, as evidenced by their lack of any non-verbal response.
4.4. Reactions to Negative System Feedback
4.4.1. Negative feedback reaction rate
We analyzed a time window starting from the system feedback announcement and lasting up to 5 seconds post-announcement, yielding a duration of 10 seconds for the first and 13 seconds for the second announcement. Of the 20 participants, 18 reacted to the negative feedback. Notably, the two non-respondents did not engage with the robot throughout. Additionally, two participants responded to only one of the two feedback statements. Out of 40 negative feedback messages, 8 prompted a verbal reaction and 34 elicited a non-verbal response. Participants exhibited a range of non-verbal reactions; while some smiled and laughed, others gestured in frustration or hesitated, seemingly confused. Eye-rolling was also observed in response to the system announcement. Verbal reactions further highlighted this confusion. One participant remarked, ”I … okay” (P18), while others expressed frustration and justified their own performance, such as ”I’m doing as much as I can do, system announcer” (P12) or the team’s performance; ”I mean we’re fast enough” (P7).
4.4.2. Blame attribution
Overall, 10 participants blamed no one for the system announcement, 4 blamed the robot, 3 the team, and 3 themselves. Among those blaming the robot, only one cited its small talk, while the rest attributed the announcements to its slow speed. One participant who blamed the team stated, “…Maybe in terms of like the team chemistry, that was probably where it was lacking”(P16). Another explained how they thought the novelty of the conversation impacted the team’s performance, “I also get a little more distracted, because …this is just …a new thing …it’s something that I guess you would get used to if you’re working with it every day”(P10).
Those who blamed themselves acknowledged possible mistakes. Meanwhile, 15 participants discussed they were confused at what prompted the system announcements because they thought the team was performing efficiently. P1 explained how the announcement “felt like out of my control. So I was like, I don’t think there was anything else that could have been done.”
Figure 4. Examples of positive and negative user interaction with the robot small talk behaviors. The top row shows a positive example where the participant expressed genuine interest in learning more about the robot (P7). The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
Figure 4: Examples of positive and negative user interaction with the robot small talk behaviors. The figure has two rows. The top row contains seven images alternating between images focusing on the participant and images focusing on the robot, with snippets of the conversation shown below indicating participant interest in conversation about the robot’s background and specifications. The bottom row contains six images alternating between images focusing on the participant and images focusing on the robot, with snippets of the conversation shown below indicating participant indicating surprise or confusion on why the robot is engaging in conversation unrelated to the task at hand.
Figure 4. Examples of positive and negative user interaction with the robot small talk behaviors. The top row shows a positive example where the participant expressed genuine interest in learning more about the robot (P7). The bottom row is an example of a user who perceived the interaction negatively and exhibited a potential mental model mismatch of expectations (P12).
5. Discussion
We explored the incorporation of conversational elements such as small talk into industrial robots, especially those not inherently built for social interactions.
While opinions on the robot’s conversational abilities vary, participants showed a willingness to engage in off-topic conversations initiated by the robot. In this section, we discuss our findings and offer design guidelines for robot small talk.
5.1. Robot-Initiated Small Talk: First Impression
Participants’ immediate impressions and engagements in small talk with the robot validate considering such dialogue for non-social robots. Our findings (section 4.3) suggest the majority of participants ($N=17$) positively received the robot’s conversation attempts, attesting to the potential of robots to engage meaningfully in small talk right from the outset. Yet, not every response was unequivocally positive. Some participants abstained from engagement ($N=2$) or posed a rhetorical question ($N=1$, see Figure 4 bottom). Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (dennler2023design, 12).
Participants’ task engagement revealed insights into their receptivity to unexpected robot behavior (small talk) for a otherwise non-social robot (section 4.3). While most participants ($N=13$) briefly glanced at the robot, suggesting a brief attention shift without disruption, the ”freeze” reactions of some ($N=4$) signaled a task disruption. This echos prior work noting that extended mutual eye-gaze in human-human exchanges often signals ”problematic interruptions” (degutyte2021role, 11). The robot’s initial conversation also evoked affective reactions. The smiles from many participants ($N=8$) suggest robots can spark positive emotions early on. Although a smile might also indicate pleasant ”surprise”, overt signs of surprise from others ($N=6$) emphasize the unexpectedness of small talk from a primarily task-oriented robot. Such shift in user attention and emotional state underscores the unexpectedness of the first question, likely due to the robot’s unforeseen conversational approach and a perceived mismatch with its physical embodiment.
In summary, the notable engagement and minimal task disruption despite surprise indicate that even if participants are surprised by the robot’s first question, it does not hinder their task. This emphasizes participants’ readiness to converse with robots, even during ongoing tasks and when the unexpected conversation is most pronounced. As participants adapt to the robot’s conversational style, diversions due to small talk might reduce further. Future research should investigate optimal small talk design to balance task efficiency and user enjoyment. Calibrating conversation frequency is vital to avoid minimize excessive talkativeness that distracts, or causes awkwardness due to prolonged silence.
5.2. Potential of Small Talk for Otherwise Non-Social Robots
5.2.1. User engagement: Small talk in purely functional tasks
While engaging in small talk with a robot arm manipulator during collaborative tasks, we observed a noticeable willingness to not only respond to the robot’s direct questions but also its comments that do not warrant a user response (section 4.1). For instance, in Figure 1, the participant comments (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”). While most of the responses by the participants were single sentenced, 12% of the responses were multiple sentences showing higher and meaningful engagement (section 4.2.1). Such level of user engagement is evident. In Figure 1, the participant’s response ”(chuckle) I think it’s a pretty classic game …They even made a movie about it.” goes beyond merely answering the robot’s query; it offers supplementary information, perhaps in anticipation of the robot’s genuine interest in the game. It is as if the participant was attempting to pique the robot’s curiosity, subtly suggesting beforehand, “You might like it”.
Although the desire to ask robot questions in return varied among participants, 45% of them posed personal questions to the robot, inquiring about its preferences or capabilities (see section 4.2.2 and Figure 4). Additionally, participants perceived the small talk positively; P20 mentioned, “it was cool to have somebody to talk to”, while P16 characterized the conversation with the robot as, “more personal and …less mechanical”. Thus, our exploratory study indicates user engagement and responsiveness to robot-initiated off-topic conversations, suggesting potential for integrating small talk to enhance user engagement in human-robot collaborative tasks that involve physical manipulation.
Potential 1: Small talk may boost meaningful engagements in human-robot interactions, even with non-anthropomorphic robots in purely functional scenarios.
5.2.2. Ice-breaking through small talk: building rapport and unveiling robot capabilities
Participants demonstrated curiosity about the robot’s functions, often posing small talk questions to probe its capabilities and background. For instance, as illustrated in Figure 4 top, one participant inquired, ”Hey robot, tell me more about yourself. I’m very interested. Like, you said you’re from Germany.” Another, referencing Figure 4 bottom, sought clarity on the robot’s primary function by asking, ”Aren’t you here just to help me with these pipes?” Such questions highlight the potential of employing conversational frameworks, even for non-social robots, to build common ground and mutual understanding in human-robot teams. By structuring initial interactions around themes such as ”getting to know your teammate”, robots can ease introductions, reducing potential barriers to seamless collaboration. Participant P7’s sentiment, “I’m curious to learn more about it” encapsulates this curiosity. Thus, small talk can serve as an initial icebreaker, fostering team unity and rapport. Over time, or in subsequent sessions, these conversations can become instrumental in communicating the robot’s capabilities, ensuring that users develops a comprehensive understanding of its functions. Furthermore, by incorporating task-related insights, tidbits, or safety tips into these dialogues, the robot can harness users’ social attention to support task performance.
Potential 2: Small talk can be an effective ice-breaker, fostering rapport and understandings of robot capabilities in collaborative settings.
5.3. Challenges and Design Guidelines for Robot Small Talk Framework
5.3.1. Dynamic and user-aware conversational framework
The conversational framework we employed for robot small talk demonstrated constraints. The design was rigid, causing interactions to feel scripted rather than organic as pointed out by P16: “the robots question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant.” The robot’s knowledge bank was confined to a limited scope of topics (by design). For instance, in the conversation with P15, after the robot’s question “By the way, do you like to travel?” , the participant responded in negation and requested next pipe: “No. Bring me a yellow pipe”. The next robot question (“What is the coolest place you have ever been to?”) was not well received by the participant. Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) by the participant. Similarly, P18 responded to the robot question about traveling: “You know I don’t know and I don’t feel like talking about traveling right now.” Thus, rigid conversational frameworks fail to adapt to the evolving nature of human conversation, highlighting the future need for a more dynamic, expansive and user-aware conversational approach.
To increase conversational flexibility, small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics. Leveraging Large Language Models (LLMs) can help tailor small talk for robots and expand a robot’s knowledge. However, they must be adapted to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences such as, ”How were you created?”, or ”Have you traveled?” Thus, LLMs can enhance the conversational capabilities but should not be utilized without necessary modifications via prompt engineering or fine tuning to avoid confusion about a robot’s physical capabilities and realistic experiences.
Design Guideline 1: Robots should use dynamic, user-aware conversational frameworks supporting flexible small talk interactions beyond rigid topics or formats.
5.3.2. Balancing small talk with task performance
Engaging in small talk with people during physical tasks poses challenges, such as with the frequency of questions posed. For instance, initiating small talk during every user command or robot action might be excessive as expressed by a participant (P15) through a rhetorical question to the robot: “Why do you talk so much?” Over-engagement in non-task related conversations can be distracting for users (veletsianos2012learners, 51, 31), potentially jeopardizing the overall performance or efficiency of the collaboration. Therefore, it is imperative to strike a balance between fostering rapport and ensuring task performance. Small talk frequency and complexity should be adapted based on task intricacy and observed team performance. During critical tasks, or when user confusion or distraction arises, it might be wise to limit or halt small talk. Using error and confusion detection frameworks (as proposed in (stiber2023using, 47, 27)) can help robots gauge progress and performance. Using these frameworks, robots can determine the appropriate time, or whether at all, to engage in small talk.
Design Guideline 2: Robots should ensure small talk is context-sensitive, enhancing team rapport without impeding task completion.
5.3.3. Mental model mismatch: managing user expectations
Several participants were unaware of interactive possibilities with the robot as evident in section 4.2.2. Participant P12 comments and inquires, “That’s strange coming from a robot. Aren’t you here to just help me with these pipes?” (see Figure 4 bottom). This points towards a potential misalignment in their mental model of the robot’s capabilities. P3’s statement, ”From beginning to end, I always felt like it’s a robot [non-human form]”, further illustrates this misalignment. It raises a pivotal question: What conversational depth is appropriate for a robot manipulator that lacks human-like form and operates within constrained physical and cognitive boundaries?
Conversely, when robots participate in conversations, it might inflate people’s expectations beyond the robot’s actual capabilities. For instance, few participants expected the robot to be more capable, only to find the robot lacking, as P7 pointed out: “I mean, because …right now, artificial intelligence is so popular. I thought a robot [is] gonna be more intelligent.”
Such misalignment may be exacerbated by participants’ familiarity with advanced conversational AI tools, such as Alexa and ChatGPT, leading them to expect that the “social” robot possesses enhanced conversational abilities. Therefore, when a physical robot starts chatting, it may lead to misalignment in user expectations and robot capabilities, as P10 expressed: “Because once [robot] starts asking me stuff, then I was …oh, maybe it’s …kind of like GPT-esque, where …has further knowledge rather than just questions, but …then my expectation went back a little bit.” If the robot’s dialogue suggests a high degree of intelligence, any shortfall in meeting those expectations could lead to increased user disappointment. Even slight mismatches between the voice and appearance of robots have been shown to introduce artifacts in human-robot interactions (mcginn2019can, 29). Establishing common ground and rapport crucially relies on instilling an accurate mental model of the robot in individuals’ minds (kiesler2005fostering, 19). Thus, it is essential to consider how user expectations can be managed adequately while designing social characteristics for robots. The robot’s social cues may shape individuals’ perceptions; if misaligned initially, the robot’s early behaviors can correct these views and recalibrate common ground (kiesler2005fostering, 19).
Design Guideline 3: Intentionally calibrate user expectations by matching robot conversational cues with its capabilities, considering users’ familiarity with AI.
5.4. Limitations and Future Work
Despite its insights, our study’s short duration and single-session focus may not fully capture the depth, intricacies and dynamics of extended human-robot collaborations and how they are influenced by small talk. Additionally, the study did not consider cultural nuances of participants, which are vital as different cultures perceive and value small talk differently. While this work serves as an exploratory examination of the use of small talk in human-robot interactions, future studies should focus on testing hypotheses to quantify the benefits of small talk for rapport building and to examine its role in human-robot collaborations with negative feedback, team evaluations, or robot errors. When designing a conversational framework for robot small talk, it is essential to research and ensure the appropriateness of content, timing, and context. For instance, the framework should be sensitive to user preferences on topics, adapt according to the user’s state for ideal timing, and consider team performance metrics to prevent disruptions during critical tasks. Lastly, future research should explore methods to align a robot’s conversational capabilities with its functional abilities. It is important to keep users apprised of the robot’s true abilities, possibly using small talk as an informative medium. Nevertheless, the strategic introduction of small talk can be instrumental in building and strengthening rapport between users and task-oriented robots in collaborative scenarios.
Acknowledgements.
This work was supported by National Science Foundation award #2141335.
Appendix
Table 3. Overview of Codes and Their Definitions for the Data Analysis
Response to First Robot Question: Conversation Reciprocation
Name Definition
Answer
Response (Yes) and answered robot’s question
Refuse
Response (Yes) but did not answer the robot’s question
No engagement
Participant did not answer the robot’s question
Response to First Robot Question: Impact on Task
Name Definition
Minimal (glance)
Looks at the robot but no task disruption
Freeze
Stops current action (minimum of 2 seconds) and looks at the robot before continuing the task
None
Did not stop or look at the robot
Response to First Robot Question: Affective Reactions
Name Definition
Pleasant
Individual exhibits a smile
Surprise
Individual exhibits behaviors such as a chin raise, lower lip raise, brow lowering, etc.
Indifference
No reaction
Non-Verbal Reactions to Negative Feedback
Name Definition
Verbal Response
An audible or verbal phrase spoken by the participant
Non-Verbal Response
Behaviors such as smiling, laughing, eye-rolls, head nods, arm movements, etc.
Blame Attribution
Name Definition
Self
Selects themselves as a potential cause for the negative feedback announcement
Robot
Selects the robot as a potential cause for the negative feedback announcement
Team
Selects the team as a potential cause for the negative feedback announcement
Nobody
Believes neither they nor the robot caused the negative feedback announcement
Motives for Initiating Questions
Name Definition
Curiosity
Individual describes a level of curiosity
Felt Human-like
Individual makes connections to human-likeness
Social Convention
Individual describes factors relating to social conventions
l6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">